Circular No. 8727

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF SATURN Further to IAUC 8523, S. S. Sheppard, D. C. Jewitt, and J. Kleyna

report the discovery of nine new satellites of Saturn (S/2004 S 19 and S/2006 S 1–S/2006 S 8) in data obtained during 2004 Dec.–2005 Mar. and 2006 Jan.–Apr. with the Subaru 8.2-m reflector, aided by B. G. Marsden’s ongoing linkages and search ephemerides. The astrometry and orbital data appear on MPECs 2006-M44, 2006-M45, and 2006-M48, showing retrograde orbits with orbital periods ranging from 862 to 1300 days.

SUPERNOVAE 2006dj, 2006dk, 2006dl Three additional supernovae have been discovered on unfiltered CCD

images: 2006dj and 2006dl by M. Baek and W. Li (LOSS/KAIT; cf. IAUC 8725), and 2006dk by M. Migliardi (communicated by A. Dimai; cf. IAUC 7373).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006dj June 5.46 22^{h}59^{m}22.^{s}89 +53^{o}44′09′′.8 18.9 0′′.5 E, 15′′.0 S
2006dk June 25.88 12 11 34.35 +57 44 26.3 16.3 6′′ E, 11′′ N
2006dl June 29.24 13 05 16.73 +25 57 27.8 18.0 9′′.5 E, 0′′.2 N

Additional KAIT magnitudes for 2006dj in UGC 12287: 2000 Oct. 7.24 UT, [20.0; 2006 June 25.46, 18.9; 26.43, 18.9. Additional CROSS magnitudes for 2006dk in NGC 4161: Jan. 30, [18.5; June 26.87, 16.2. Nothing appears at the location of 2006dk on Palomar Sky Survey infrared, red, and blue plates. A spectrogram obtained on June 26.92 shows that SN 2006dk is a type-II supernova close to the time of explosion (details on CBET 565). Additional magnitudes for 2006dl in MCG +04-31-5: June 19.26, [19.5; 30.23, 18.0.

COMETS C/2006 J12 AND C/2006 K6 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8724; GS = G. Sun) — both described by K. Battams as being small and stellar in appearance, with no tails. C/2006 K6 reached mag 6.5 on May 17.246 UT at 5.5R_{⊙} in C2 images.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 J12 May 15.646 3^{h}31.^{m}6 +17^{o}23′ C2  GS 2006-M28
C/2006 K6  17.071 3 38.1 +17 24 C3/2 RM 2006-M28
2006 June 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green